OpenStudy (anonymous):

this is all like another language hahaha...solve the quadratic equation by completing the square?? 3x^2-18x+24=0???

OpenStudy (anonymous):

make your life easier by factoring out a 3 and then writing \[3 (x-4) (x-2) = 0\]

OpenStudy (anonymous):

okay well i got the factor by 3 but dont you have to include and the 3x^2 somehow thats how i kept screwing up?

OpenStudy (anonymous):

you want to solve this equation, so if you can factor it is easier than using the formula.

OpenStudy (anonymous):

now you know \[x-4=0\] so \[x=4\] or \[x-2=0\] so \[x=2\]

OpenStudy (anonymous):

ohhh gotcha!!